1. You own the following stocks in your portfolio. What is the beta of your portfolio....

1.

You own the following stocks in your portfolio. What is the beta of your portfolio.

Stock Invested Amount Beta
A 6,249 1.39
B 8,522 1.07
C 7,420 1.98
D 4,140

1.43

2. You purchased one share of Kohl's Corporation, Inc for $47.10 per share. The company paid a dividend of $5.10 per share during the year, and had an ending share price of $49.38. What is the capital gains yield?
Answer #1

1.Total investment=(6249+8522+7420+4140)=26331

Portfolio beta=Respective beta*Respective weight

=(6249/26331*1.39)+(8522/26331*1.07)+(7420/26331*1.98)+(4140/26331*1.43)

=1.46(Approx).

2.Capital gains yield=(End value-Beginning value)/Beginning value

=(49.38-47.1)/47.1

=4.84%(Approx).